There are 850 douglas fir trees and ponderosa pine trees in a section of forest bought by sawz logging co. The company paid an average of $300 for each douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy? McGraw Hill- Beginning Algebra, Chapter 2 exercise 6 and 7 for both problems. NO ISBN Sorry!
Thankyou so very much!!!!!!!

Answer by funmath(2873)

Firs+ponderosas=850
Let fir=x
Then the rest of the 850 trees are ponderosas.
Let ponderosa=850-x
Cost of firs=300x
Cost of ponderosas=225(850-x)
Total:217,000
The equation you have to solve:
300x+225(850-x)=217,000
300x+191,250-225x=217,000
(300-225)x+191,250=217,000
75x+191,250=217,000
75x+191,250-191,250=217,000-191,250
75x=26,250
75x/75=26,250/75
x=350
Fir trees:x=350
Ponderosa pines:850-x=850-350=500